# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v1

@@ Abstract
Large Language Model (LLM) agents are reshaping the game industry, particularly with more intelligent and human-preferable game characters. However, existing game benchmarks fall short of practical needs: they lack evaluations of diverse LLM capabilities across various game genres, studies of agentic modules crucial for complex gameplay, and fine-tuning datasets for aligning pre-trained LLMs into gaming agents. To fill these gaps, we present Orak, a foundational benchmark designed to train and evaluate LLM agents across diverse real-world video games. Unlike existing benchmarks, Orak includes 12 popular video games spanning all major genres, enabling comprehensive studies of LLM capabilities and agentic modules essential for intricate game scenarios. To support consistent evaluation of LLMs, we introduce a plug-and-play interface based on Model Context Protocol (MCP) that enables LLMs to seamlessly connect with games and manipulate agentic modules. Additionally, we propose a fine-tuning dataset, consisting of LLM gameplay trajectories across diverse game genres. Orak offers a comprehensive evaluation framework, encompassing general game score leaderboards, LLM battle arenas, and in-depth analyses of visual input state, agentic strategies, and fine-tuning effects, establishing a foundation towards building generic gaming agents. Code is available at https://github.com/krafton-ai/Orak.
@@ 1 Introduction
Large Language Model (LLM) agents are revolutionizing various industries [1], and well-established benchmarks play a key role in advancing their capabilities on complex tasks, e.g., coding [2, 3, 4], web search [5, 6, 7], and scientific research [8, 9]. Similarly, in the game industry, there is growing interest in leveraging LLM agents to enhance users’ game experiences, e.g., introducing more intelligent non-player characters (NPCs), monsters, and companions [10]. In response, several benchmarks have been proposed to assess the capability of LLMs in playing games [11].
@@
While these benchmarks have effectively utilized games to evaluate LLMs’ general capabilities, they exhibit three major limitations: 1) they mostly rely on text-only games or 2D-grid simulators rather than complex real video games, 2) they offer insufficient assessment of agentic modules, such as self-reflection, memory, and tool use, which are essential to complex gameplay, and 3) they lack fine-tuning datasets necessary to adapt pre-trained LLMs into effective gameplay agents, which significantly hinder the adoption of LLM agents in real-world video games.
@@
To this end, we present Orak, a foundational benchmark designed to evaluate LLM agents across diverse video games. As shown in Figure 1, Orak includes 12 video games played by millions to billions of users worldwide: Street Fighter III, Super Mario, Ace Attorney, Her Story, Pokémon Red, Darkest Dungeon, Minecraft, Stardew Valley, StarCraft II, Slay the Spire, Baba Is You, and 2048. These games span 6 major game genres, i.e., action, adventure, role-playing, simulation, strategy, and puzzle, enabling a comprehensive assessment of key abilities required for general gameplay; action games enable testing fine-grained player control, adventure games challenge long-term memory and error handling, and strategy/puzzle games require complex logical reasoning and multi-step planning. In addition, with the use of real video games, Orak ensures evaluation on rich, dynamic environments with varying stages, levels, and story-driven quests, which are even challenging for humans.
@@
To enable consistent evaluation of rapidly evolving LLMs, we introduce a plug-and-play interface for interaction with game environments and agentic modules using Model Context Protocol (MCP) [12]. Each game environment and agentic module package operates as an independent MCP server, providing game mechanics (e.g., retrieving game states, executing game steps) or agentic strategies (e.g., reflection, planning) as callable functions to LLMs. During gameplay evaluation, the LLM interacts with these servers by sequentially retrieving game states, performing action inference using agentic modules, and executing game steps. This interface enables streamlined evaluation across diverse games and supports controlled studies of various agentic modules.
@@
Furthermore, we propose a fine-tuning dataset that aims to transform pre-trained LLMs into effective gaming agents. The dataset consists of game interaction trajectories generated by expert LLMs, e.g., GPT-4o, using various agentic strategies on all games in Orak. These trajectories encapsulate meta-knowledge on how and when to use the agentic strategies to play various genres of games, leading to more resource-efficient and effective gaming LLM agents.
@@
Our benchmark offers comprehensive evaluation dimensions, including general game scores with leaderboards, competitive LLM battle arenas, and in-depth analyses of visual input state, agentic strategies, and fine-tuning effects. Extensive experiments on Orak with 12 LLMs reveal that (1) proprietary LLMs generally achieve superior performance across games, with significant gaps compared to open-source LLMs, (2) their performance gap becomes narrow in battle games, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) visual states often hinder gameplay performance, and (5) fine-tuning on gameplay trajectories enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game and out-of-distribution (OOD) game scenarios. We believe that Orak not only establishes a foundation for developing gaming LLM agents but also serves as a critical benchmark for evaluating general LLMs on realistic, long-horizon decision-making tasks.
@@
Figure 1: Overview of Orak, a benchmark designed to evaluate LLM agents across 12 real-world video games from various genres. Using Model Context Protocol (MCP) as a plug-and-play interface, it ensures efficient and reliable assessment, supporting gameplay leaderboards, battle arenas, and in-depth studies on agentic modules and fine-tuning.
@@ 2 Related Work
2.1 Playing Games by LLMs
@@
Many recent works have explored the use of LLMs for gameplay. Early efforts focused on text-based games such as Jericho [13], Zork [14], and TextCraft [15], where LLMs leveraged their abilities in exploration and reasoning to navigate textual environments. Subsequent work shifted toward 2D-grid games, including Chess [16], NetHack [17], and Crafter [18], where spatial reasoning and puzzle-solving skills became more important for successful gameplay. More recently, several studies have applied LLMs, combined with agentic workflows, to play more complex video games. Notably, LLMs have been used for item crafting in Minecraft [19, 20], city building in Civilization [21], Pokémon battles [22], and strategic planning in StarCraft [23]. However, these approaches rely on manually designing agentic workflows customized for each specific game given proprietary LLMs, e.g., GPT [24], limiting their usability toward developing a general gaming agent.
2.2 Evaluation Benchmarks for LLMs with Games
@@
As gameplay requires complex cognitive abilities, e.g., context understanding, logical reasoning, and error handling, several recent benchmarks have sought to evaluate LLMs or Vision Language Models (VLMs) on games [11]. GAMA-Bench [25], GameBench [26], GameArena [27], and SmartPlay [28] focus primarily on text-based games, assessing LLMs’ ability to navigate and reason within textual environments. Barlog [29] and LVLM-Playground [30] are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs. Using video games, Cradle [31] evaluates VLMs through 3 simulation games and 1 adventure game, V-MAGE [32] assesses VLMs on 5 action video games, and DSGBench [33] validates LLMs on 6 strategic games. Despite their contributions, existing benchmarks have several limitations: they lack coverage of diverse game genres, omit in-depth studies on agentic modules essential for complex gameplay, and rely solely on visual inputs, despite VLMs’ current limitations in spatial perception and reasoning [30, 32]. Also, aligning pre-trained LLMs with agentic workflows for effective gameplay remains largely unexplored. Table 1 summarizes the key characteristics of game benchmarks.
@@
Benchmarks Game Domain Full Genre ##\## Games State Type Agent Study Fine-tuning Set
GAMA-bench [25] Text 8 Text
GameBench [26] Text 9 Text
GameArena [27] Text 6 Text
SmartPlay [28] Text/2D-grid 6 Text
Balrog [29] Text/2D-grid 6 Text/Image
LVLM-Playground [30] 2D-grid 6 Image
Cradle [31] Video 4 Image
V-MAGE [32] Video 5 Image
DSGBench [33] Video 6 Text
Orak (Ours) Video 12 Text/Image
@@
Table 1: Game Benchmark Comparison. ‘Full Genre’ indicates whether six common game genres are fully covered (i.e., action, adventure, role-playing, simulation, strategy, and puzzle). ‘State Type’ refers to the modality of game state provided to LLMs or VLMs. Unlike prior benchmarks, Orak is the only benchmark that features a diverse set of real video games, fully covers all major genres, supports both text and image state inputs for LLMs/VLMs, provides ablation studies for agent modules, and releases a fine-tuning dataset.
2.3 Fine-tuning LLMs toward Agents
@@
Recent advances in LLM agents have enabled applications in diverse interactive settings, including web navigation, API usage, and game environments. To enhance agent capability, recent research efforts have proposed agentic strategies such as chain-of-thought reasoning [34, 35], self-reflection [36, 37], hierarchical task planning [38, 39, 40, 41], and self-generated skill libraries [20]. Complementing these advancements, efforts have also focused on fine-tuning strategies tailored to LLM agents. These strategies span two main directions: data-centric approaches, which involve supervised fine-tuning on curated expert demonstrations [42, 43, 44, 45, 46, 47]; and framework-oriented approaches, which focus on learning from agentic interactions [48, 49, 50, 51, 52]. Notably, FireAct [42] emphasizes unified data formatting and CodeAct [47] highlights the importance of high-quality curation of training trajectories. However, fine-tuning methods for game-playing agents remain insufficiently explored. Unlike structured tasks in web, programming, or math domains, games involve large, dynamic, and partially observable state spaces. This requires agents to generalize across a wide variety of situations and learn diverse behavior patterns, posing unique challenges.
@@ 3 Orak
Figure 2: Evaluation pipeline of Orak. Game scores are computed via eval.py by simply configuring game, LLM backbone, and agentic strategy. Orak supports two types of submissions: (1) customizing llm.py with new backbone LLMs, and (2) customizing agent.py with new agentic strategies. The agentic strategies are callable by the LLM through the MCP interface in eval.py (in grey box).
@@ 3.1 Game Environments
LLM Capabilities Required. LLM agents require various capabilities to play video games. Figure 3 summarizes the level of LLM capabilities required to play each game in Orak, which is measured on a scale of 1 to 3, following the principled criteria adopted from [28].
  • Rule Following (RF): LLM agents should be able to follow game-specific rules. The level is measured by the extent to which adherence to rules is required for gameplay (1: single rule, 2: fewer than 5 rules, 3: 5 or more rules).
  • Logical Reasoning (LR): The number of LLM’s reasoning hops required to determine an in-game action (1: 1 hop, 2: 1 to 3 hops, 3: 3 or more hops).
  • Spatial Reasoning (SR): The level of spatial understanding required for gameplay (1: not necessary, 2: required in specific situations, 3: critical to core gameplay).
  • Long-text Understanding (LTU): The extent of long-context comprehension required for gameplay (1: a few lines, 2: a few paragraphs, 3: longer than one page with 500$+$ words).
  • Long-term Planning (LP): The extent to which strategic planning is required (1: not necessary, 2: planning for up to 3 sequential actions, 3: essential to plan more than 3 sequential actions).
  • Error Handling (EH): The extent to which error correction is required during gameplay (1: not necessary, 2: requires a one-step rollback, 3: requires multi-step rollback and re-planning).
  • Odds Handling (OH): The extent to which understanding randomness is required for gameplay (1: not necessary, 2: randomness exists in game, 3: randomness is critical to core gameplay).
The level for each game is measured by 8 human participants, and the moderate value is reported. Since most video games are designed to require various cognitive abilities even for humans, they tend to require high levels of LLM capabilities in many dimensions. For example, action games, Street Fighter III and Super Mario in red, require spatial reasoning and rule following, more than long-context understanding and planning, while adventure games, Ace Attorney and Her Story in yellow, emphasize long-text understanding and logical reasoning due to the need to comprehend long storylines. More detailed analysis of required LLM capabilities for each game is in Appendix B.
@@
(a) Street Fighter III [53] is a 2D side-scrolling fighting action game with 20 unique characters, each equipped with distinctive skills. (1) Game state: The player character, opponent character, remaining time, each player’s health, super-bar, stun-bar gauges, and the distance between the two characters. (2) Action space: 15-20 discrete actions: ‘move closer’, ‘low punch’, ‘high kick’, etc. (3) Evaluation task: Beating the game bot; performance is measured by the number of stages cleared.
@@
(b) Super Mario [54] is a side-scrolling game where the player controls Mario to avoid obstacles, defeat enemies, and reach the flag. (1) Game state: The positions (x,y) and sizes of obstacles and enemies extracted from the current game state. (2) Action space: Mario keeps moving to the right, and LLM decides the jump level, discretized in 6 bins. (3) Evaluation task: Reaching out to the final flagpole; performance is measured by the horizontal distance that Mario travels before dying.
@@
(c) Ace Attorney [55] is a courtroom adventure game where players act as defense attorneys, gathering evidence and cross-examining witnesses to prove their client’s innocence. (1) Game state: Dialogue history, collected evidence, court records profiles, etc. (2) Action space: Player’s courtroom actions: advancing dialogue, accessing court records, pressing witnesses, and presenting evidence. (3) Evaluation task: Performance is measured by response correctness and total steps taken.
@@
(d) Her Story [56] is an interactive adventure game where players explore police interview clips to uncover a hidden truth. (1) Game state: History of queries and search results with metadata for the first 5 clips (visual description, date, viewing status, and transcript if played). (2) Action space: Searching for clips with keywords, or selecting a video to play. (3) Evaluation task: Uncover the truth; Performance is measured by the number of distinct video clips viewed to complete the game.
@@
(e) Pokémon Red [57] is a turn-based role-playing game where a player explores, collects Pokémon, and battles other trainers to progress the storyline. (1) Game state: Player’s location, party Pokémon (species, level, HP, status), inventory, battle state, and screen text. (2) Action space: Choosing high-level tools or low-level joypad actions. (3) Evaluation task: Defeat Brock, the first gym leader; Progress measured by how many of 12 predefined storyline flags are triggered within a step limit.
@@
(f) Darkest Dungeon [58] is a turn-based role-playing game where heroes explore dungeons while managing stress and resources. (1) Game state: Party status (character stats, health, stress, and status effects), available skills, and enemy encounters. (2) Action space: Combat actions like ‘attack’, ‘heal’, and ‘swap’. (3) Evaluation task: Complete the first expedition; Performance is measured by the sum of the successful combats, the survived heroes, and their remaining stress capacities.
@@
(g) Minecraft [59] is an open-ended sandbox game where players explore a world, gather resources, and survive by placing and breaking blocks. (1) Game state: The player’s position, inventory, health status, the nearby blocks and biome, etc. (2) Action space: Executable JavaScript code within the Mineflayer environment [60]. (3) Evaluation task: Crafting a target item; performance is measured by whether the item is collected in the inventory.
@@
(h) Stardew Valley [61] is an open-ended life simulation game where players farm, fish, mine, and explore. (1) Game state: Player’s location, energy, inventory, crop status, soil status, date, and weather. (2) Action space: Leaving the house, entering the house, sleeping, buying seeds, tilling soil, watering, harvesting, and selling crops. (3) Evaluation task: Earning the most money by harvesting crops within the first 13 in-game days; performance is measured by total profit.
@@
(i) StarCraft II [62] is a real-time strategy game where players gather resources, construct buildings, train units, and command armies to defeat opponents. (1) Game state: Resource levels, unit/building counts, production queues, research progress, and observed enemy info. (2) Action space: 72 discrete actions, including unit training, building, research, and strategic operations. (3) Evaluation task: Beating built-in AI bots; performance is measured by the win rate.
@@
(j) Slay the Spire [63] is a deck-building roguelike game where players ascend a multi-floor tower, battling enemies and building decks. (1) Game state: Player’s class, deck, hand, health, relics, energy, enemies’ intents and statuses, and current floor. (2) Action space: Playing a card during combat, ending the turn, and selecting a card reward after combat. (3) Evaluation task: Defeating the final boss at the top floor; performance is measured by the number of floors reached.
@@
(k) Baba Is You [64] is a puzzle game where a player manipulates the rules by moving word tiles on a board. (1) Game state: Coordinates of text and object tiles, and active rules. (2) Action space: A single movement ‘up’, ‘down’, ‘left’, and ‘right’, or a sequence of such moves. (3) Evaluation task: Solving the first stage; if the stage is not cleared, partial credit is awarded based on sub-goals (e.g., breaking the ‘Wall Is Stop’ rule).
@@
(l) 2048 [65] is a sliding tile puzzle game that aims to combine numbered tiles on a 4$\times$4 grid board to create a tile of the value 2048. (1) Game state: The current configuration of the 4$\times$4 grid, where each cell contains either a number (power of 2) or is empty. (2) Action space: Four discrete actions; ‘up’, ‘down’, ‘left’, and ‘right’. (3) Evaluation task: Creating the 2048 tile; performance is measured by the normalized progress toward creating the 2048 tile.

### v2

@@ Abstract
Large Language Model (LLM) agents are reshaping the game industry, by enabling more intelligent and human-preferable characters. Yet, current game benchmarks fall short of practical needs: they lack evaluations of diverse LLM capabilities across various game genres, studies of agentic modules crucial for complex gameplay, and fine-tuning datasets to adapt pre-trained LLMs into gaming agents. To fill these gaps, we present Orak, a benchmark for training and evaluating LLM agents across 12 popular video games spanning all major genres. Using a plug-and-play interface built on Model Context Protocol (MCP), Orak supports systematic and reproducible studies of agentic modules in varied game scenarios. We further release a fine-tuning dataset of expert LLM gameplay trajectories covering multiple genres, turning general LLMs into effective game agents. Orak offers a united evaluation framework, including game leaderboards, LLM battle arenas, and in-depth analyses of input modality, agentic strategies, and fine-tuning effects, establishing a foundation towards versatile gaming agents. Code is available at https://github.com/krafton-ai/Orak.
@@ 1 Introduction
Large Language Model (LLM) agents are revolutionizing various industries (Wang et al., 2024a), and well-established benchmarks play a key role in unlocking their abilities on complex tasks, e.g., Coding (Hendrycks et al., 2021, Zhuo et al., 2024), Web Search (Liu et al., 2023, Pan et al., 2024, Levy et al., 2024), and Scientific Research (Mühlbacher et al., 2024, Zhang et al., 2025). In game industry, there is growing interest in using LLM agents to enhance user game experiences e.g., more intelligent non-player characters (NPCs), monsters, or companions (NVIDIA, 2025). In response, many benchmarks have been proposed to assess the ability of LLMs to play games (Hu et al., 2024b).
@@
While these benchmarks have effectively utilized games to evaluate LLMs’ general capabilities, they exhibit three major limitations: 1) they often rely on text-only games or 2D-grid simulators rather than complex real video games, 2) they offer insufficient assessment of agentic modules, such as self-reflection, memory, and tool use, which are essential to complex gameplay, and 3) they lack fine-tuning datasets necessary to adapt pre-trained LLMs into effective gameplay agents, which significantly hinder the adoption of LLM agents in real-world video games.
@@
To this end, we present Orak, a foundational benchmark designed to evaluate LLM agents across diverse video games. As shown in Figure 1, Orak includes 12 video games played by millions to billions of users worldwide: Street Fighter III, Super Mario, Ace Attorney, Her Story, Pokémon Red, Darkest Dungeon, Minecraft, Stardew Valley, StarCraft II, Slay the Spire, Baba Is You, and 2048. These games span 6 major game genres, i.e., action, adventure, role-playing, simulation, strategy, and puzzle, enabling a comprehensive assessment of key abilities required for versatile gameplay; action games enable testing fine-grained player control, adventure games challenge long-term memory and error handling, and strategy/puzzle games require complex logical reasoning and multi-step planning. With the use of real video games, Orak also ensures evaluation on rich, dynamic environments with varying stages, levels, and story-driven quests, which are even challenging for humans.
@@
To enable consistent evaluation of rapidly evolving LLMs, we introduce a plug-and-play interface using Model Context Protocol (MCP) (Hou et al., 2025), allowing LLMs to seamlessly interact with agentic modules in gameplay. Each game environment and agentic module package operates as an independent MCP server, providing game mechanics (e.g., retrieving game states, executing game steps) or agentic strategies (e.g., reflection, planning) as callable tools to LLMs. During evaluation, LLM interacts with these servers by sequentially retrieving game states, performing action inference using agentic modules, and executing game steps, which enables streamlined evaluation across diverse games and supports controlled studies of various agentic modules.
@@
In addition, we release a fine-tuning dataset that aims to transform pre-trained LLMs into effective gaming agents. The dataset consists of game interaction trajectories generated by expert LLMs, e.g., GPT-4o, using core agentic strategies on all games in Orak. These trajectories encapsulate meta-knowledge on how and when to use the agentic strategies to play various genres of games, leading to more resource-efficient and effective gaming agents.
@@
Our benchmark offers comprehensive evaluation dimensions, including game score leaderboards, competitive LLM battle arenas, and in-depth analyses of visual input state, agentic strategies, and fine-tuning effects. Extensive experiments on Orak with 15 LLMs reveal that (1) proprietary LLMs achieve superior performance across games with significant gaps from open-source LLMs, (2) their performance gap becomes narrow in battle scenarios, (3) proprietary LLMs benefit from extended agentic workflows, while open-source LLMs show limited gains, (4) visual states often hinder gameplay performance yet, and (5) fine-tuning enables effective transfer of gameplay meta-knowledge from larger LLMs to smaller ones, leading to generalization in intra-game, out-of-distribution (OOD) game, and non-game unseen scenarios like math and web interaction. We believe that Orak not only establishes a foundation for building gaming agents but also serves as a critical benchmark for evaluating general LLMs on realistic, long-horizon decision-making tasks.
@@
Figure 1: Overview of Orak, a benchmark designed to train and evaluate LLM agents across 12 video games across genres. Using MCP as a plug-and-play interface, it ensures systematic assessment, supporting gameplay leaderboards, battle arenas, and studies on agentic modules and fine-tuning.
@@ 2 Related Work
Playing Games by LLMs.
@@
Many works have explored the use of LLMs for gameplay. Early efforts focused on text-based games such as Jericho (Hausknecht et al., 2020), Zork (Tsai et al., 2023), and TextCraft (Prasad et al., 2023), where LLMs navigate textual environments with reasoning. Subsequent work shifted toward 2D-grid games, including Chess (Feng et al., 2023), NetHack (Küttler et al., 2020), and Crafter (Hafner, 2021), where spatial reasoning and puzzle-solving skills became more important. More recently, several studies have applied LLMs, combined with agentic workflows, to play more complex video games, such as Minecraft (Fan et al., 2022, Wang et al., 2023), Civilization (Qi et al., 2024), Pokémon (Hu et al., 2024c), and StarCraft (Ma et al., 2024). However, these approaches rely on manually customized agentic workflows for each specific game, limiting their usability toward developing a general gaming agent.
Evaluation Benchmarks for LLMs with Games.
@@
As gameplay requires complex cognitive abilities, e.g., context understanding, logical reasoning, and error handling, several recent benchmarks have sought to evaluate LLMs or Vision Language Models (VLMs) on games (Hu et al., 2024b). GAMA-Bench (Huang et al., 2024), GameBench (Costarelli et al., 2024), GameArena (Hu et al., 2024a), and SmartPlay (Wu et al., 2023) focus on text-based games, assessing LLMs’ ability to navigate and reason on textual environments. Barlog (Paglieri et al., 2024) and LVLM-Playground (Wang et al., 2025) are mainly based on 2D-grid games, such as TicTacToe and Chess, to evaluate the spatial and visual reasoning capabilities of LLMs/VLMs. Using video games, Cradle (Tan et al., 2024) evaluates VLMs in 1 adventure and 3 simulation games, V-MAGE (Zheng et al., 2025) assesses VLMs on 5 action games, DSGBench (Tang et al., 2025) validates LLMs on 6 strategic games, and LMGame-Bench (Hu et al., 2025) mainly focus on puzzle games. Despite their contributions, existing benchmarks lack full coverage of game genres, omit in-depth studies on agentic modules, often rely on visual inputs, and under-explore the way to align pre-trained LLMs into versatile game agents. Table 1 summarizes the key characteristics of game benchmarks.
@@
Benchmarks Game Domain Full Genre #\# Games Model Type Agent Ablation Fine-tuning Set
GAMA-bench (Huang et al., 2024) Text 8 LLM
GameBench (Costarelli et al., 2024) Text 9 LLM
GameArena (Hu et al., 2024a) Text 6 LLM
SmartPlay (Wu et al., 2023) Text/2D-grid 6 LLM
Balrog (Paglieri et al., 2024) Text/2D-grid 6 LLM/VLM
LVLM-Playground (Wang et al., 2025) 2D-grid 6 VLM
Cradle (Tan et al., 2024) Video 4 VLM
V-MAGE (Zheng et al., 2025) Video 5 VLM
DSGBench (Tang et al., 2025) Video 6 LLM
LMGame-Bench (Hu et al., 2025) Video 6 LLM/VLM
Orak (Ours) Video 12 LLM/VLM
@@
Table 1: Game Benchmark Comparison. ‘Full Genre’ means whether six major genres are fully covered. ‘Model Type’ indicates whether the benchmark supports LLMs or VLMs. Unlike prior benchmarks, Orak is the only benchmark that fully covers all major genres, supports both LLMs/VLMs, provides ablation studies for agent modules, and releases a fine-tuning set.
Fine-tuning LLMs toward Agents.
@@
To enhance agent capability, many efforts have proposed agentic strategies, e.g., chain-of-thought reasoning (Wei et al., 2022, Yao et al., 2023), self-reflection (Shinn et al., 2023, Park et al., 2023), hierarchical task planning (Huang et al., 2022b; a), and skill libraries (Wang et al., 2023). Complementing these advancements, efforts have focused on fine-tuning strategies tailored to LLM agents, which span two main directions: data-centric approaches, which involve fine-tuning on curated expert demonstrations (Chen et al., 2023, Zeng et al., 2023, Chen et al., 2024); and framework-oriented approaches, which focus on learning from agentic interactions (Feng et al., 2024, Chen et al., 2025, Putta et al., 2024). Notably, FireAct (Chen et al., 2023) emphasizes unified data formatting and CodeAct (Wang et al., 2024b) highlights the importance of high-quality curation of training trajectories. However, fine-tuning methods for game-playing agents remain largely explored. Unlike structured tasks in web, programming, or math domains, games involve large, dynamic, and partially observable state spaces, which requires agents to generalize across a variety of situations and learn diverse behavior patterns, posing unique challenges.
@@ 3 Orak
Figure 2: Evaluation pipeline of Orak. Game scores are computed via eval.py by simply configuring game, LLM backbone, and agentic strategy. Orak supports two types of submissions: (1) customizing llm.py with new backbone LLMs, and (2) customizing agent.py with new agentic strategies. The agentic strategies are callable by LLMs via MCP interface in eval.py (in grey box).
@@ 3.1 Game Environments
LLM Capabilities Required.
LLM agents require diverse capabilities to play games. Figure 3 summarizes the capability levels needed for each game in Orak, measured using principled criteria adopted from the context of game design (Wu et al., 2023, Koster, 2013):
  • Rule Following (RF). The level is measured by the extent to which adherence to rules is required for gameplay (1: single rule, 2: fewer than 5 rules, 3: 5 or more rules).
  • Logical Reasoning (LR). The number of LLM’s reasoning hops required to determine an in-game action (1: 1 hop, 2: 1 to 3 hops, 3: 3 or more hops).
  • Spatial Reasoning (SR). The level of spatial understanding required for gameplay (1: not necessary, 2: required in specific situations, 3: critical to core gameplay).
  • Long-text Understanding (LTU). The extent of long-context comprehension required for gameplay (1: a few lines, 2: a few paragraphs, 3: longer than one page with 500++ words).
  • Long-term Planning (LP). The extent to which strategic planning is required (1: not necessary, 2: planning for up to 3 sequential actions, 3: essential to plan more than 3 sequential actions).
  • Error Handling (EH). The extent to which error correction is required during gameplay (1: not necessary, 2: requires a one-step rollback, 3: requires multi-step rollback and re-planning).
  • Odds Handling (OH). The extent to which understanding randomness is required for gameplay (1: not necessary, 2: randomness exists in game, 3: randomness is critical to core gameplay).
@@
The capability level is measured on a scale of 1 to 3 by 8 human participants, and the moderate value is reported. Since most video games are designed to require various cognitive abilities even for humans, they tend to require high levels of LLM capabilities in many dimensions. For example, action games, Street Fighter III and Super Mario in red, require spatial reasoning and rule following, more than long-context understanding and planning, while adventure games, Ace Attorney and Her Story in yellow, emphasize long-text understanding and logical reasoning due to the need to comprehend long storylines. More detailed analysis of required LLM capabilities for each game is in Appendix B.
@@
(a) Street Fighter III (Capcom, 1997) is a 2D side-scrolling fighting action game with 20 unique characters, each with distinctive skills. (1) Game state: The player character, opponent character, remaining time, each player’s health, super-bar, stun-bar gauges, and the distance between the two characters. (2) Action space: 15-20 discrete actions: ‘move closer’, ‘low punch’, ‘high kick’, etc. (3) Evaluation task: Beating the game bot; performance is measured by the number of stages cleared.
@@
(b) Super Mario (Christian Kauten, 2018) is a side-scrolling game where the player controls Mario to avoid obstacles, defeat enemies, and reach the flag. (1) Game state: The positions (x,y) and sizes of obstacles and enemies extracted from the current game state. (2) Action space: Mario keeps moving to the right, and LLM decides the jump level, discretized in 6 bins. (3) Evaluation task: Reaching out to the final flag; performance is measured by the horizontal distance that Mario travels before dying.
@@
(c) Ace Attorney (Capcom, 2001) is a courtroom adventure game where players act as defense attorneys, gathering evidence and witnesses to prove their client’s innocence. (1) Game state: Dialogue history, collected evidence, court records profiles, etc. (2) Action space: Player’s courtroom actions: advancing dialogue, accessing court records, pressing witnesses, and presenting evidence. (3) Evaluation task: Performance is measured by response correctness and total steps taken.
@@
(d) Her Story (Barlow, 2015) is an interactive adventure game where players explore police interview clips to uncover a hidden truth. (1) Game state: History of queries and search results with metadata for the first 5 clips (visual description, date, viewing status, and transcript if played). (2) Action space: Searching for clips with keywords, or selecting a video to play. (3) Evaluation task: Uncover the truth; Performance is measured by the number of distinct video clips viewed to complete the game.
@@
(e) Pokémon Red (Game Freak, 1996) is a role-playing game where a player explores, collects Pokémon, and battles other trainers to progress the storyline. (1) Game state: Player’s location, party Pokémon (species, level, HP, status), inventory, battle state, and screen text. (2) Action space: Choosing high-level tools or low-level joypad actions. (3) Evaluation task: Defeat Brock, the first gym leader; Progress measured by how many of 12 predefined storyline flags are triggered.
@@
(f) Darkest Dungeon (Red Hook Studios, 2016) is a role-playing game where heroes explore dungeons while managing stress and resources. (1) Game state: Party status (character stats, health, stress, and status effects), available skills, and enemy encounters. (2) Action space: Combat actions like ‘attack’, ‘heal’, and ‘swap’. (3) Evaluation task: Complete the first expedition; Performance is measured by the sum of the successful combats, the survived heroes, and their remaining stress.
@@
(g) Minecraft (Mojang Studios, 2011) is an open-ended sandbox game where players explore a world, gather resources, and survive by placing and breaking blocks. (1) Game state: The player’s position, inventory, health status, the nearby blocks and biome, etc. (2) Action space: Executable JavaScript code within the Mineflayer environment (contributors, 2013). (3) Evaluation task: Crafting a target item; performance is measured by whether the item is collected in the inventory.
@@
(h) Stardew Valley (ConcernedApe, 2016) is an open-ended life simulation game where players farm, fish, mine, and explore. (1) Game state: Player’s location, energy, inventory, crop status, soil status, date, and weather. (2) Action space: Leaving the house, entering the house, sleeping, buying seeds, tilling soil, watering, harvesting, and selling crops. (3) Evaluation task: Earning the most money by harvesting crops within the first 13 in-game days; performance is measured by total profit.
@@
(i) StarCraft II (Blizzard Entertainment, 2010) is a real-time strategy game where players gather resources, construct buildings, train units, and command armies to defeat opponents. (1) Game state: Resource levels, unit/building counts, production queues, research progress, and observed enemy info. (2) Action space: 72 discrete actions, including unit training, building, research, and strategic operations. (3) Evaluation task: Beating built-in AI bots; performance is measured by the win rate.
@@
(j) Slay the Spire (MegaCrit, 2017) is a deck-building roguelike game where players ascend a multi-floor tower, battling enemies and building decks. (1) Game state: Player’s class, deck, hand, health, relics, energy, enemies’ intents and statuses, and current floor. (2) Action space: Playing a card during combat, ending the turn, and selecting a card reward after combat. (3) Evaluation task: Defeating the final boss at the top floor; performance is measured by the number of floors reached.
@@
(k) Baba Is You (Hempuli, 2019) is a puzzle game where a player manipulates the rules by moving word tiles on a board. (1) Game state: Coordinates of text and object tiles, and active rules. (2) Action space: A single movement ‘up’, ‘down’, ‘left’, and ‘right’, or a sequence of such moves. (3) Evaluation task: Solving the first stage; if the stage is not cleared, partial credit is awarded based on sub-goals (e.g., breaking the ‘Wall Is Stop’ rule).
@@
(l) 2048 (Cirulli, 2014) is a sliding tile puzzle game that aims to combine numbered tiles on a $\times4$ grid board to create a tile of the value 2048. (1) Game state: The current configuration of the $\times4$ grid, where each cell contains either a number (power of 2) or is empty. (2) Action space: Four discrete actions; ‘up’, ‘down’, ‘left’, and ‘right’. (3) Evaluation task: Creating the 2048 tile; performance is measured by the normalized progress toward creating the 2048 tile.